the length of a rectangle is 5 ft more than twice its width, and the area of the rectangle is 88ft squared. Find dimensions
Answer by rapaljer(4671)

Let x = the width
2x+5= the length

The equation is A=L*W
x(2x+5)=88
2x^2+5x-88=0

I hope this factors:
(2x-11)(x+8)=0

There appear to be two solutions:
2x-11=0 or x+8=0
2x=11 or x=-8 This answer is rejected since a side of a rectangle can't be negative.
x=11/2 or 5.5 ft = width

2x+5=Length
2(11/2)+5=11+5=16 = Length

Check: Area = L*W = 16*5.5 = 88sq ft.
It checks.

Now, if you need help with a number of topics in this problem, please see own my website by clicking on my tutor name "Rapaljer" anywhere in algebra.com. Then click on the link there to my website. From my Homepage, look for the link "Basic, Intermediate, and College Algebra: One Step at a Time." If you need help with Solving Quadratic Equations, then see "Basic Algebra" and look fo this topic in Chapter 2. If you need help factoring this trinomial, then look in "Basic Algebra", then "Chapter 2" for the topic "Advanced Trinomial Factoring." This is FULL of examples, exercises, and answers to all the exercises. See also the "Math in Living Color" pages where many of these exercises are solved IN COLOR. My own students used to tell me that my explanations are much easier to understand than the traditional textbooks. Perhaps you will like this section. I just want to make math easier for people who have trouble with math. It's usually not as hard as everyone thinks it is. The website has a lot of resources--all FREE.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus